# Analyzing Logs of Large-Scale Software Systems using Time Curves Visualization

### v1

@@ IV-D Multiple curve analysis
The logging output of software systems can be widely different. However, two or more instances of the same application may be running concurrently, thus increasing the likelihood of their output being similarly structured. In this section, we use our approach to overlay the projection curves of multiple application instances to identify outliers or branching differences.
Figure 4: Multiple curve analysis of three instances of the same application. The systems are shown in different color and style, namely solid purple, dashed yellow and dot-and-dashed green. Labels A to E represent stages of interest.
To illustrate a multiple curve analysis, we gathered log collections corresponding to three instances of a simple server application that receives connections from clients. The application was monitored from launch to shutdown, with the main processing stage being listening to remote connections. Figure 4 shows the Time Curves of the three log collections overlaid on top of each other. Note that each system is represented with a different color (and line style) to ease visualization. In addition, due to the fact that different colors are used to represent different systems, the time dimension is no longer represented besides the order of checkpoints. The labels A to E represent the following states:
  • Label A corresponds to the startup sequence. It can be observed that all three curves begin at nearly the exact location, due to the initialization being identical except the small variations, caused by e.g., different execution paths or process PIDs. These cause a very small distance that enables to differentiate between exactly equal executions and semantically equal executions.
  • The startup continues until label B where the three systems enter the processing stage. The applications are listening to remote connections. Notice that each one receives a variable number of connections yet they are visually clustered in label B.
  • Until stage C, processing had remained equal for all applications. However, it can be observed that two of the curves, namely the yellow dashed and the dot-and-dashed green, enter an outlier state in label C before returning to the normal processing (label B). In fact, by examining the logs corresponding to label C it can be seen that the two curves threw a An illegal reflective access operation has occurred followed by additional error logs. These do not occur in the solid purple curve.
  • After normal processing has resumed, all applications commence the shutdown sequence from label D to E.
@@ IV-E Scalability of the approach
In this section, we show how the method scales both in terms of size and complexity of the input data. By size we refer to number of input log lines (which approximately results in a similar number of log records), whereas for complexity we mean the distribution of the lengths of the records and the number of extracted templates. On one hand, and as explained in Section III-G, longer log lines have an impact on performance due to the intrinsic quadratic term of the Levenshtein distance. On the other hand, more templates also result in a performance penalty because of the pairwise template comparison between checkpoints. While this can be become a potential bottleneck, further optimizations may be applied to improve performance as a trade off for accuracy, such as using a computationally cheaper string metric as replacement of the Levenshtein distance (e.g., the q-gram distance [30]) or limiting the number of extracted templates during the clustering stage.
The experiments described in this section were run 10 times and averaged. Further details in the hardware setup can be found in [22].
Figure 5: Average runtime and throughput as a function of the number of log messages of the Hadoop file system logs. Note that x-axis is in $\log_{2}$ scale. The left y-axis represents average runtime in seconds using stacked bars, whereas the second y-axis represents average throughput using a dashed line. Stacked bars are decomposed proportionally by runtime.
Figure 5 shows the average runtime and throughput in terms of processed records per second for an increasing number of logs. The proportion of time required by each of the stages is represented as part of the stacked bars. Note that the dataset source is fixed to maintain the distribution of the log records. As can be observed, peak performance is achieved with the largest dataset containing four million log records at a processing rate of roughly 35,000 log records per second (with a single core). The smaller datasets do not contain enough log lines to pay off for (1) constant initialization factors, such as importing of libraries, (2) multidimensional projection, which depends on number of checkpoints and goodness of fit rather than number of log records, (3) curve annotation, etc. These penalties get smoothed out as the actual core processing (grouping, template extraction and distance computation) takes over when the number of records increases. It can be concluded that the runtime (and throughput) of the method is linear at worst given that the x-axis is in logarithm scale.
Figure 6: Average runtime and throughput as a function of the dataset. The left y-axis represents average runtime in seconds using stacked bars, whereas the second y-axis represents average throughput using a dashed line. Stacked bars are decomposed proportionally by runtime.
@@ V Conclusions and future work
In this manuscript, we have shown that our proposal is able to summarize and represent large collections of log messages while scaling efficiently and providing insights. This is achieved via the following contributions:
@@
    An automated grouping methodology for detecting the main events of a collection of logs based on cumulative information (such as timestamps and severity level) without requiring prior knowledge.
@@
    Several extensions to the Time Curves visualization technique, including multiple curve comparison, enriching events with logging information or animating curves.
@@
    The complete experimental approach, serving as a reference for validating analysis of log datasets, which can be reused and compared with future methods.
Furthermore, our results show that the approach is in fact able to explain the logs from three different stream processing frameworks and how the similarity projection reveals further properties of the experiment (such as the resemblance between startup and failure recovery in Kafka Streams in Section IV-A) which had not been foreseen. In the same line, we have also analyzed a public dataset and then contrasted it both manually and with our automated approach.
@@
It should be noted that in this paper, we have focused on analyzing complete log datasets. However, working with subsets according to some criteria is also of interest. For example, splicing datasets by a desired property (e.g., a certain component of an application such as the ExecutionGraph in Apache Flink) can produce more focused and tailored insights.
In conclusion, we believe the proposed approach offers a holistic application that can ease up system log analysis, thus resulting in a significant time reduction for a variety of scenarios. This includes, among others, analyzing system-wide problems and performance bottlenecks, identifying security breaches and threats, troubleshooting and debugging application errors, monitoring and auditing user activity, analyzing resource management, etc.
In further research, we plan to address:
@@
    Speed up processing of even larger datasets by exploiting the parallelism potential, such as in the preprocessing, grouping and clustering stages.
@@
    The event detection mechanism may benefit from exploring the use of (1) Fast Fourier Transform for faster convolutions, (2) adaptive thresholds to refine the boundary between begin and end of events and (3) additional information such as sources of tags, keywords, performance metrics, etc.

### v2

@@ IV-D Multiple curve analysis
The logging output of software systems can widely differ. However, some instances of the same application may be running concurrently, thus increasing the likelihood of their output being similarly structured. In this section, we use our approach to overlay the projection curves of multiple application instances to identify outliers or branching differences.
Figure 5: Multiple curve analysis of three instances of the same application. Each system is represented with a different color and line style to ease visualization. In addition, due to the fact that different colors are used to represent different systems, the time dimension is no longer represented besides the order of checkpoints. Labels A to E represent stages of interest.
To illustrate a multiple curve analysis, we gathered log collections corresponding to three instances of a simple server application that receives connections from clients. The application was monitored from launch to shutdown, with the main processing stage being listening to remote connections. Figure 5 shows the Time Curves of the three log collections overlaid on top of each other. Labels A to E represent the following stages:
  • Label A corresponds to the startup stage. It can be observed that all three curves begin at nearly the exact location, due to the initialization being identical except the small variations, caused by e.g., different execution paths or process IDs. These cause a very small distance that enables to differentiate between exactly equal executions and semantically equal executions.
  • The startup continues until B where the three systems enter the processing stage. The applications are listening to remote connections. Notice that each one receives a variable number of connections yet they are visually clustered in B.
  • Until stage C, processing had remained equal for all applications. However, it can be observed that two of the curves, namely the yellow dashed and the dot-and-dashed green, enter an outlier checkpoint in C before returning to the normal processing (Label B). In fact, by examining the logs corresponding to C it can be seen that the two curves threw a "An illegal reflective access operation has occurred" followed by error logs, which do not occur in the solid purple curve.
  • After normal processing has resumed, all applications begin the shutdown stage from D to E.
@@ IV-E Scalability of the approach
As Dynatrace deployments keep growing, the efficiency of log analytics becomes extremely important, especially given the scale of the log ingest. Therefore, in this section, we show how the method scales both in terms of size and complexity of the input data. By size we refer to number of input log lines (which approximately results in a similar number of log records), whereas for complexity we mean the distribution of the lengths of the records and the number of extracted templates. On one hand, longer log lines have an impact on performance due to the intrinsic quadratic term of the Levenshtein distance. On the other hand, more templates also result in a performance penalty because of the pairwise template comparison between checkpoints. While this can be become a potential bottleneck, further optimizations may be applied to improve performance as a trade off for accuracy, such as using a computationally cheaper string metric as replacement of the Levenshtein distance (e.g., the q-gram distance [30]) or limiting the number of extracted templates during the clustering stage.
All experiments described in this section were averaged over 10 runs. Further setup details can be found in [8].
Figure 6: Average runtime and throughput as a function of the number of log messages of the Hadoop file system logs. Note that x-axis is in $\log_{2}$ scale. The left y-axis represents average runtime in seconds using stacked bars, whereas the second y-axis represents average throughput using a dashed line. The proportion of time required by each of the stages is represented as part of the stacked bars.
Figure 6 shows the average runtime and throughput in terms of processed records per second for an increasing number of logs. Note that the dataset source is fixed to maintain the distribution of the log records. As can be observed, peak performance is achieved with the largest dataset containing four million log records at a processing rate of roughly 35,000 log records per second (with a single core). The smaller datasets do not contain enough log lines to pay off for (1) constant initialization factors, such as importing of libraries, (2) multidimensional projection, which depends on number of checkpoints and goodness of fit rather than number of log records, (3) curve annotation, etc. These penalties get smoothed out as the actual core processing (grouping, template extraction and distance computation) takes over when the number of records increases. It can be concluded that the runtime (and throughput) of the method is linear at worst given that the x-axis is in logarithm scale.
Figure 7: Average runtime and throughput as a function of the dataset. The left y-axis represents average runtime in seconds using stacked bars, whereas the second y-axis represents average throughput using a dashed line. The proportion of time required by each of the stages is represented as part of the stacked bars.
@@ V Conclusions and future work
In this manuscript, we have shown that our proposal is able to summarize and represent large collections of log messages while scaling efficiently and providing meaningful insights. This is achieved via the following contributions:
@@
    An automated grouping methodology for detecting events in a collection of logs based on cumulative information (such as timestamps and severity level) without requiring any prior knowledge.
@@
    Several extensions to the Time Curves visualization technique, including multiple curve comparison, enriching data points with logging information or animations.
@@
    The experimental approach, serving as a reference for validating analysis of log datasets, which can be reused and compared with future methods.
Furthermore, our results show that the approach is in fact able to explain the logs collected from three different stream processing frameworks and how the similarity projection reveals further properties of the framework logging which had not been foreseen (such as the resemblance between startup and failure recovery in Kafka Streams in Section IV-A). In the same line, we have also analyzed a public dataset and then compared our manual explanations with the automated results from the proposed method.
@@
It should be noted that in this paper, we have focused on analyzing complete log datasets. However, working with subsets according to some criteria (splicing datasets by a desired property) can theoretically produce more focused and tailored insights, e.g. analyzing only a certain component of an application such as the ExecutionGraph in Apache Flink.
In conclusion, we believe the proposed approach offers a holistic method that can ease up system log analysis, thus speeding up analysis of system-wide problems and performance bottlenecks, identification of security threats, etc.
In further research, we plan to address the following:
@@
    Speed up processing of even larger datasets by exploiting the parallelism potential, such as in the preprocessing, event detection or clustering.
@@
    The event detection mechanism may benefit from exploring the use of (1) Fast Fourier Transform for faster convolutions, (2) adaptive thresholds to refine the boundary between begin and end of events and (3) additional information such as performance metrics.